OpenStudy (anonymous):

If a gallon of gas costs 334 9 /10 cents, how much does 9 gallons cost?

OpenStudy (anonymous):

334 9/10 = 3349/10 = 334.90 now just multiply 334.90 x 9 and you will get your answer Do you understand ?

OpenStudy (anonymous):

yess, thank you

OpenStudy (anonymous):

your welcome :)

OpenStudy (anonymous):

i got 3014.1 and i guess its wrong...

OpenStudy (anonymous):

Thats what I got....$ 3014.10

OpenStudy (anonymous):

Just think about it...if it is 334.90 per gallon, 9 gallons is gonna cost quite a bit

OpenStudy (anonymous):

yeah..thank you

OpenStudy (anonymous):

not a problem...thats what I am here for :)

OpenStudy (anonymous):

(: